Problem 3: Migrating Snowy Owls (16 points) Snowy Owls are an Artic species that are sometimes observed at Northeastern Marine Center on Nahant in the winter. Bird banding data has measured their migration distances (in km) and found the following data for a sample of 20 owls: ## 317 323 379 398 442 456 517 539 563 573 584 609 646 731 735 770 801 833 944 956 Calculate the mean, standard deviation, standard error and 95% confidence intervals on migra- tion distance. Show your work by setting up a table where each row is an observation, the first column is the data and the second colum is the squared deviation from the mean. 4 points a)

B) Draw an approximate boxplot for the data, where the whiskers of the boxplot extend the range of the data. Your y-axis should span from 0 to 1200.


c) Draw a barplot on the same y-axis as you made for the boxplot. For the barplot, the height of the bar should be mean. If sampling was repeated many times for a sample size of 20 from the population and the mean was recalculated, approximately what percent of sampled means would fall within the error bars on the boxplot? 4 points the mean, and the error bar should be 1 standard error above and below the d) Compare the boxplot and the barplot. What kind of different information does each convey about the data? Discuss the distribution of the data, its skewness, the location and dispersion of the data, and confidence on the mean. 4 points
Answer #1

a)

deviation from mean Data 317 323 379 398 442 456 517 539 563 573 584 609 646 731 735 770 801 833 944 956 Mean - 605.8 83405.4 79975.8 51438.2 43180.8 6830.4 22440.0 7885.4 4462.2 1831.8 1075.8 475.2 10.2 1616.0 15675.0 16692.6 26961.6 38103.0 51619.8 114379.2 122640.0 Total Sq. Dev.-710699.2 Variance37405.2

Here, mean is computed using the formula:

317 323 -944-956 605.s Mean

And Squared Deviation from Mean, (eg. for 1st observation)

Squared Deviation = (317 - 605.8)2 = 83405.4

Summing the squared deviations and dividing the sum by ( No. of observation - 1) = (20-1) = 19, we obtain the variance.

(323-605.8 (956-605, 8)2 + (317-605.8 Variance 19

  10699.205 = 37405.2 19

And Standard deviation is nothing but the square root of the variance.

Standard deviation = V37405.2- 193.4

Standard error is computed using the formula:

Standard error = Standard deviation / sqrt{n}

  1919 44.37

b)

1200.0 1000.00 800.00 600.00 400.00 200.00 0.00

where the upper whisker corresponds to the value 956 and the lower whisker to the value 317.

Bar plot:

1孑 500 千OD 300 200

d) Comparing boxplot and barplot, the box plot reveals the median, quartiles and the range of the values in the whole dataset, whereas in bar plot we have plotted the mean and its range of error of 1 SE.

The data appears to be slightly positively skewed.The 95% confidence interval for mean is given by:

605.8 2.58 * 44.37 (491.3.720.3)

The mean is supposed to lie within this interval for about 99% of the time.